OpenStudy (anonymous):

Consider the line that passes through the point (3, -6) and has a slope of 4. Write the equation in point slope form.

jimthompson5910 (jim_thompson5910):

y - y1 = m(x - x1) ... point slope form y - y1 = 4(x - x1) ... plug in the given slope m = 4 y - (-6) = 4(x - 3) ... plug in the given point (3, -6) y + 6 = 4(x - 3) ... simplify and that's all there is to it

OpenStudy (anonymous):

y= 4x-18

jimthompson5910 (jim_thompson5910):

we stop there because they want it in point slope form, not slope intercept form

OpenStudy (anonymous):

Thanks :)